OpenStudy (anonymous):

Help asap please! c: Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3).

OpenStudy (anonymous):

y = 2x + 2 for a perpendicular line, the slope is the negative reciprocal. So the slope is the negative reciprocal of 2 which is - 1/2. y = mx + b ....using slope -1/2 and points (6,3) 3 = -1/2(6) + b 3 = - 3 + b 3 + 3 = b 6 = b your equation is : y = - 1/2x + 6 (slope intercept form) or if you need it in standard form : Ax + By = C, then it is ... y = - 1/2x + 6 (multiply everything by 2) 2y = - x + 12 x + 2y = 12 <--standard form

OpenStudy (anonymous):

Thank You! (:

OpenStudy (anonymous):

not a problem...glad I could help :)